OpenStudy (anonymous):

William is a contractor. He is remodeling his home and has decided to do most of the work himself. Occasionally his family will help him with some of the work. William wants to start replacing the molding on his dining room ceiling and window. William measures the dining room ceiling to see how much molding he will need to go around the perimeter of the room's ceiling. William measured the length of the dining room to be 8 3/4 ft and the width to be 10 7/8 ft. How much molding does William need for the dining room ceiling? Show your work.

OpenStudy (danjs):

The perimeter of the room will be Perimeter = 2*Length + 2*Width

OpenStudy (anonymous):

So... That means?

OpenStudy (danjs):

add those 4 sides, that gives you the perimeter of the room

OpenStudy (danjs):

Perimeter of rectangular room is 2 times the length + 2 times the width. I put the mixed numbers into improper fractions to make it easier to add them.

OpenStudy (danjs):

\[Perimeter = 2*[\frac{ 27 }{ 4 } *\frac{ 2 }{ 2 }]+ 2*\frac{ 87 }{ 8 }\] The fraction over 4 is multiplied by 2/2 to give the common denominator of 8 so you can add them together.

OpenStudy (danjs):

\[Perimeter = \frac{ 108 }{ 8 } + \frac{ 174 }{ 8 } = \frac{ 282 }{ 8 }~~ft\]

OpenStudy (danjs):

reduce the fraction , or turn into a mixed number if you need to .

OpenStudy (anonymous):

Would that be 141/4?

OpenStudy (danjs):

yes, or 35 1/4 ft.

OpenStudy (anonymous):

Okay :D!

OpenStudy (danjs):

hey , one second, there is a miscalculation

OpenStudy (danjs):

I messed up turning one of the numbers into improper fractin... here is the correct one...

OpenStudy (danjs):

\[Perimeter = 2*\frac{ 35 }{ 4 }*\frac{ 2 }{ 2 } + 2*\frac{ 87 }{ 8 }\]

OpenStudy (danjs):

\[Perimeter = \frac{ 140 }{ 8 } + \frac{ 174 }{ 8 }\] \[Perimeter = \frac{ 314 }{ 8 }\]

OpenStudy (danjs):

or 157/4 which is 39 1/4 ft

OpenStudy (danjs):

sorry bout that.

OpenStudy (danjs):

If you wanted to do it w/o converting mixed numbers to improper fractions, you take 2 of each of the lengths.... 2*(10) + 2*(7/8) + 2*(8) + 2*(3/4)

OpenStudy (danjs):

20 + 14/8 + 16 + 6/4 36 + 14/8 + 12/8 36 + 26/8 36 + 13/4 36 + 3 + 1/4 39 1/4